OpenStudy (anonymous):

In the equation 2x + 2y = 5, the y-intercept is 5 true or false?

OpenStudy (inkyvoyd):

do you know how to rewrite an equation into slope-intercept form?

OpenStudy (inkyvoyd):

ok so basically slope itnercept form looks like y=mx+b where m and b are numbers so if you have an equation that's 3x+2y=-4 to write it into slope intercept form you just solve for y so 2y=-3x-4 y=-3/2x-2 the nice thing about slope intercept form is that once you have something in slope intercept form that number without the x behind it (the "b" of y=mx+b) is your y intercept

OpenStudy (inkyvoyd):

so let's try it with your equation. you have 2x+2y=5 so you want to isolate y so it's the only thing, right?

pooja195 (pooja195):

Well thats a little complicated why not just sub the X for 0

OpenStudy (inkyvoyd):

@pooja195 there's a number of ways to do it, but I have a feeling that her curriculum will want her to use the slope intercept form more than the definition of y intercept

OpenStudy (anonymous):

for some reason will not let me clear messages and type back ok im ready taking notes on this stuff

OpenStudy (inkyvoyd):

ok so let's start with your equation. 2x+2y=5

OpenStudy (inkyvoyd):

so we want to get y alone on one side. Any ideas how to do it?

OpenStudy (inkyvoyd):

ok so let's subtract 2 x from both sides... then the 2x on the left side cancels out and goes to the right side 2x+2y=5 -2x =-2x 0x+2y=5-2x right?

OpenStudy (inkyvoyd):

so 0x equals 0 so we just have 2y=5-2x right?

OpenStudy (inkyvoyd):

now all we have to do is get rid of the 2 on both sides. we want y insteasd of 2y, so we just divide 2y by 2. But we have to do that to the other side of the equation too

OpenStudy (inkyvoyd):

so 2y/2=(5-2x)/2 we have to do each individual term here y=5/2-2x/2 then we simplify and put x first so it's in the form y=mx+b y=-x+5/2

OpenStudy (anonymous):

ok im writing stuff down and following you on this one

OpenStudy (inkyvoyd):

so does everything so far make sense?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

so we're actually really close to our answer here. because we've written our equation in the form y=mx+b, all we need to do is identify b so our equation is y=-x+5/2 can you identify b?